the half-life of a radioactive isotope is the time it takes for a quantity of the isotope to be reduced to half its initial mass. starting with 165 grams of a radioactive isotope, how much will be left after 3 half-lives?
Answer by Theo(13342)

i believe it will be 165 * (1/2)^3 which will be equal to 165 * 1/8 which will be equal to 20.625.

1/2 of 165 = 82.5
1/2 of 82.5 = 41.25
1/2 of 41.25 = 20.625